# A Survey on Brain-inspired Deep Learning via Predictive Coding

### v2

@@ 5 Plausibility of Predictive Coding in Neuroscience
In standard PC networks, synaptic values, which are crucial for neural computation, are generally not subject to strict limitations after an update. This can cause the values to become highly positive or negative, which makes the model less stable overall. This is tackled in NGC, which introduces a constraint to ensure that the Euclidean norm of any row or column in the synaptic matrix does not exceed one, aiding in stability, a practice rooted in early classical sparse coding linear generative models [76]. Beyond this, there have been recent attempts to establish heuristic limits on the magnitude of synaptic values in order to thwart declines in classification performance [154]. A second implausibility of PC models is the frequent alterations in the signs of their synapses (or “sign-flipping”), that can go from negative to positive (and vice versa) during training, an aspect that is pivotal in emulating real cortical functionality. A potential solution to prevent sign-flipping would be to enforce non-negativity constraints on synaptic values and explicitly model groups of excitatory and inhibitory neurons, given that this configuration would apply the necessary positive and negative forces essential for PC inference and learning [155, 156, 157].
6 On Software Frameworks
Software plays a crucial role in engendering innovation, design, and development of models and algorithms for computational intelligence. In the realm of machine learning, software libraries provide the necessary back-end support and optimization needed for the GPU-driven, multi-CPU computation that powers the simulation of deep neural networks. Frameworks such as Theano [158], Caffe [159], Torch/PyTorch [160], and TensorFlow [161] have been among the key software platforms that have proven instrumental, both historically as well as for modern-day efforts in machine learning and neurobiological modeling. Furthermore, software libraries focused on particular subclasses of neural network and machine learning models facilitate reproducibility, an important element of scientific practice.
With respect to PC, software frameworks that facilitate research in PC are relatively sparse. Most research in PC design and development results in disjoint, scattered model and paper-specific code and scripts, often making it difficult to generalize those models and algorithms beyond the context in which they were developed. This prohibits the wider adoption of the principles that underlie PC and accompanying variational procedures; further preventing the education of future generations of researchers that could make important contributions to PC.
Figure 7: The theory-software-hardware cycle of predictive processing. In this setting, it is shown that theory and mathematical/computational development informs the structure and functionality of the required simulation software. The relationship between software and hardware has notably been depicted to be naturally bi-directional: software can inform the design of the hardware platform and vice-versa, i.e., co-design.
Nevertheless, there do exist several software libraries that could democratize PC research. Notably, ngc-learn [162], a general simulation and development library for computational neuroscience models, was one of the earliest historical libraries to support the construction of arbitrary PC models. Its theoretical grounding in neuronal cable theory and compartmental neurons allows computational modelers to prototype and simulate neural circuits, including those based on NGC and BC-DIM. To perform deep learning experiments, PCX [122] is a newer library built on top of Equinox, a deep learning framework in JAX, and hence allows an easy plug-and-play of deep learning models, trained using the neuronal updates of PC. Furthermore, the paper [122] accompanying the library also presents state-of-the-art results in the field for a large number of frameworks. Other examples of PC libraries include pypc [163], a library for the training of supervised and unsupervised hierarchical PC models, and predify [164], a PC-like software library that allows for the conversion of various deep learning architectures to systems that incorporate elements of PC, such as its top-down predictive circuitry. Newer PC-specialized libraries still emerge to this day, offering new elements of functionality such as specialized support for stochastic and Bayesian formulations of PC as in pyhgf [165]. Fig. 7 depicts the general PC, software, and hardware interaction cycle.
7 On Novel Hardware
Hardware limitations have historically played a pivotal role in shaping the direction of research by either enabling or hindering the practical implementation of theoretical ideas [166]. At present, graphics processing units (GPUs) and tensor processing units (TPUs) are predominantly used for training deep neural networks. However, emerging technologies such as memristors, spintronics, and optics have the potential to revolutionize the landscape [167, 168, 169]. Adapting algorithms to benefit from these new hardware technologies is crucial for advancing the field. PC is particularly well-suited for the development of algorithms for alternative hardware, as it is a simple energy based model that can be trained using equilibrium propagation, and bi-level optimization [15, 170]. It is also able to conduct iterative inference with layer-wise parallel computation, Hebbian synaptic adaptation, and optimisation resting on local (variational free) energy functions. These properties (in common with most neuro-mimetic schemes) could enable a parallelization of its underlying operations, thus allowing for significant efficiency gains in training large-scale neural networks. The parallel nature of PC reduces communication and wait time, which are bottlenecks in the traditional backprop algorithm. Furthermore, the alignment of PC with neuromorphic hardware creates an opportunity for enabling energy-efficient learning methodologies [171]. A promising direction is that of using dynamical memristors, perfectly suitable for developing brain-inspired neuromorphic systems, which offer both high energy efficiency as well as high computing capacity [172].
Biophysical Implementation.
Moving beyond conventional hardware, PC is poised to play a crucial role in “intelligence-in-a-dish” technology [173, 174, 175]. This involves adaptive neuromorphic computation through non-silicon bio-physical mediums. Due to the close association of PC with cortical computation and structure [176, 134, 177], it could be inherently compatible with organoid intelligence for performing fundamental calculations [178]. The early attempts to apply PC at the spike-communication level [179, 155] are particularly promising to mimic predictive processing in a dynamic context. Furthermore, PC is well-aligned with the constraints of biophysically oriented computations. Not only would these computational frameworks be natural first candidates for emulation in biophysical substrates, but their generalization ability might vastly improve given the natural evolution that the organoid structure lends itself. That is, the organoid’s “growth” and “decay” might offer a powerful means of facilitating model selection [180], improving the long-term generalization ability of a neural system. The road to robust synthetic general intelligence and neuromorphic cognition might lie in the intersection of neuroscientific (computational) models, such as PC, and the bio-engineering of neuronal micro-physiological systems.
7.1 Spiking Predictive Coding
A potential challenge for PC (and other neurobiological credit assignment processes in machine learning) is generalizing to spike-level temporal processing. Biological neurons communicate through discrete signals known as action potentials, creating spike trains that are sparse yet contain rich timing information [181]. These spike trains serve as the inspiration for spiking neural networks (SNNs), so-called third-generation neural networks, which encode information through the precise timing of spikes [182]. This sparse communication inherent to SNNs is one of their core strengths, allowing them to consume less energy through the use of high-information content spike trains, facilitating the construction of energy-efficient hardware to support their processing, i.e., neuromorphic hardware [183, 184, 185], where deep ANNs, on the other hand, require the use of energy-intensive, top-of-the-line graphics processing units (GPUs) for effective training. Unfortunately, with a few exceptions, most modern incarnations of PC fail to account for how learning and inference could occur under such a sparse, discrete communication scheme. This may reflect the fact that PC is, under the hood, a variational scheme that reproduces the idealized behavior of a sampling scheme (e.g., particle filtering) based on single particles or neurons: cf., [124]. That is, PC, by construction, describes ensemble or population dynamics, not the behavior of individual particles or neurons. For a comprehensive survey of spiking PC, we refer the reader to [186].
Figure 8: Spiking neural coding—a formulation of PC/NGC at the spike-train level, where $\mathbf{s}^{\ell}(t)$ represents an action potential (or spike) node, and $\mathbf{x}^{\ell}(t)$ is now treated as a trace meant to approximate a value node. Solid black arrows represent forward generative synapses, dashed black arrows represent error transmission synapses, and blue dashed arrows represent non-synaptic transformations (such as the recurrent electrical current and voltage calculations underlying a spike node $\mathbf{s}^{\ell}(t)$ or the trace/value node update underlying $\mathbf{x}^{\ell}(t)$). Note that $\mathbf{o}(t)$, which could be a static image or a video, is typically converted to a spike vector $\mathbf{s}^{0}(t)$ at time $t$ (resulting in a sensory-level Poisson spike train).
The Spiking Neural Coding Framework.
The spiking neural coding framework is an attempt to formulate how PC could be implemented in spiking networks [187]. It builds on the premise that spike-based events can induce synaptic adjustments [155]. The framework employs an event-driven local adjustment rule that leverages variable traces for pre-synaptic and post-synaptic spikes. In SNNs, synaptic efficacies are updated using spike-timing-dependent plasticity (STDP) [188], where correlations in neural activities based on spike timing are used for synaptic modulation. Adapting deep SNNs using STDP is challenging and often requires carefully hand-crafted layers. In contrast, spiking neural coding utilizes a more flexible approach. which involves synaptic conductance models, membrane potential models, spike emission functions, and trace mechanisms. Each layer in this framework attempts to estimate the trace activity of another layer, with the discrepancy being passed locally backward through iteratively evolved error feedback synapses (see Fig. 8). Importantly, spiking neural coding is adaptable to various spiking functions, from the leaky integrator to the Hodgkin-Huxley model [189]. Other implementations of spike-level PC have recently begun to emerge [190, 191, 140]. Notably, in contrast to the purely unsupervised nature of the original spiking neural coding framework, models such as [190] operate in a discriminative context. However, these newer models do not address the weight transport problem, because feedback synapses are shared with the forward/generative synapses.
@@ 8 Looking Forward: Important Directions for Predictive Coding Research
While it is often claimed that PC is the optimal inversion scheme for hierarchical Gaussian generative models, its performance can be quite below modern-day deep neural networks trained with backprop. The future of PC in machine intelligence strongly depends on our ability to address and fill this gap. Specifically, among the main directions for future work related to PC should be understanding the fundamental reasons behind this performance mismatch and using this acquired insight to develop and design novel PC schemas, mathematical frameworks, and computational models that work in large scale settings, where current deep learning models excel.
@@
Future research will also need to focus on the study of optimization techniques that have proven useful and invaluable for variational inference, such as those that promote the inclusion of precision weighting parameters into the picture [198]. Despite interesting developments obtained over the last several years, and despite the fact that precisions are of primary importance in simulations used in the neurosciences, these techniques have only been tested in small and medium scale settings [199, 200, 201]. More generally, the field of deep learning has enjoyed immense benefits from simple optimization tricks developed throughout the last decade, such as dropout [202], batch normalization [203], adaptive learning rates such as Adam [204] and RMSprop [205], and the introduction of the ReLU activation [206] and its variants. Without these techniques, emerging from the joint effort of thousands of researchers, the training of extremely overparametrized neural models would not lead to the results that we observe today. Interestingly, backpropagation itself exhibited many flaws and limitations at the beginning that precluded it scaling to high-dimensional data spaces: in the late 90’s, kernel learning methods were dominant and generally found to be more effective than artificial neural networks, which were generally disregarded due the problem of vanishing gradients [207] as well as their strong requirement for heavy computational power. Most of these problems have now been addressed in over thirty years of research, again through a collective effort of thousands of people. Given the history of deep learning, one might ask: what are the dropout, batch norm, and Adam optimizer equivalents for PC? Addressing such a question would be of vital importance if the goal is to scale the applicability of PC, and will hopefully become a more prominent topic of interest of more research efforts going forward.
@@
Another probabilistic generative modeling research direction for PC is to incorporate the uncertainty about model (i.e., synaptic) parameters into the variational bound on marginal likelihood. That is, one can equip synaptic parameters with probability distributions (as opposed to using point estimates). This move would take PC closer to its Bayesian roots: current machine learning implementations of PC do not treat the model parameters as random variables and can therefore be regarded as an expectation maximization (EM) procedure, where the M-step ignores uncertainty about the parameters [209]. The benefit of treating parameters as random variables is that one can evaluate the model evidence required for structure learning [210, 211, 212, 213], i.e., one could place prediction errors on the synaptic weights and evaluate the resulting variational free energy (or marginal likelihood) for structure learning or, in a biological setting, morphogenesis [214, 133]. This research direction takes PC into the world of generalized filtering and dynamic causal modeling, in which parameters can be regarded as latent states that change very slowly [103, 215]. In the case of dynamic causal modeling (i.e., the variational inversion of state-space models under Gaussian assumptions), structure learning can be rendered particularly efficient through particular instances of Bayesian model selection [216, 217, 218]; namely, Bayesian model reduction to remove or prune redundant parameters from a full or parent model [219]. Related generalizations of PC might also leverage advances in nonparametric Bayes to learn the structure of neuronal networks. Please see [217, 220, 221] for further details.
Crafting Cognitive Control Systems.
@@
PC brings with it the promise of learning a powerful generative process that is continuously and iteratively refined as more sensory samples are gathered over time. This has led some early work to consider such a process as the basis for world models that drive modular, brain-motivated cognitive models, capable of combining perception and action in the context of playing video games [91, 111] and robotic control tasks [114, 112, 113], as well as large-scale cognitive architectures [119, 86, 120]. This has important implications beyond machine learning, particularly for the domains of cognitive science and cognitive neuroscience where a key pathway is to craft computational theories of mind and to examine their fit to human subject data on controlled psychological tasks [222] as well as their ability to generalize to cognitive functionality. Hence, a promising future direction is to design modular, increasingly complex systems made up of PC circuitry, that rely on single fundamental PC circuits and low-level dynamics.
@@
Using PC as a fundamental neural building block for cognitive control systems is still in its earliest stages, and although more effort will be needed for it to become viable for building robust computational theories of mind, it presents a worthwhile long-term direction for crafting simulated natural intelligence. This promise is made even stronger when considering that deep backprop-based networks are beginning to find use in the field of cognitive science [226]. Making developments along this direction would also crucially benefit the research in active inference [227, 85, 19, 18], given that agent systems that engage in epistemic foraging often center around the use of dynamic generative models and can even be viewed as simple cognitive control models.

### v3

@@ 5 Plausibility of Predictive Coding in Neuroscience
In standard PC networks, synaptic values, which are crucial for neural computation, are generally not subject to strict limitations after an update. This can cause the values to become highly positive or negative, which makes the model less stable overall. This is tackled in NGC, which introduces a constraint to ensure that the Euclidean norm of any row or column in the synaptic matrix does not exceed one, aiding in stability, a practice rooted in early classical sparse coding linear generative models [163]. Beyond this, there have been recent attempts to establish heuristic limits on the magnitude of synaptic values in order to thwart declines in classification performance [164]. A second implausibility of PC models is the frequent alterations in the signs of their synapses (or “sign-flipping”), that can go from negative to positive (and vice versa) during training, an aspect that is pivotal in emulating real cortical functionality. A potential solution to prevent sign-flipping would be to enforce non-negativity constraints on synaptic values and explicitly model groups of excitatory and inhibitory neurons, given that this configuration would apply the necessary positive and negative forces essential for PC inference and learning [165, 166, 167, 168].
6 Software Frameworks and Novel Hardware
Software and hardware play crucial roles in enabling innovation and practical implementation of computational intelligence models. While frameworks like PyTorch [169] and TensorFlow [170] have been instrumental for deep learning, PC research faces unique challenges due to sparse software support and evolving hardware opportunities. In terms of software, available frameworks remain relatively sparse, with most research resulting in scattered, paper-specific code that hinders wider adoption and reproducibility. Nevertheless, several libraries aim to democratize PC research: ngc-learn [171] is the official library for NGC, is grounded in neuronal cable theory and supports arbitrary PC model and neuromorphic system construction as well as general neural circuit simulation. PCX [81] is a deep learning oriented library, is built on JAX/Equinox and enables plug-and-play deep learning with PC updates and demonstrates state-of-the-art results across frameworks. Other libraries include pypc [172] for hierarchical PC models, predify [173] for converting deep architectures to PC-like systems, and pyhgf [174] for stochastic and Bayesian PC formulations, and the recently proposed jPC [87] for fully parallel deep learning architectures.
Figure 7: The theory-software-hardware cycle of predictive processing showing bidirectional relationships between theoretical development, simulation software, and hardware platforms.
6.1 Novel Hardware Implementation
Hardware limitations historically shape research directions [175]. While GPUs and TPUs dominate current training, emerging technologies like memristors, spintronics, and optics could revolutionize the field [176, 177, 178]. PC is particularly well-suited for alternative hardware as an energy-based model trainable through equilibrium propagation [15, 179]. Its iterative inference with layer-wise parallel computation, Hebbian adaptation, and local energy optimization enables significant parallelization, reducing communication bottlenecks inherent in backpropagation. This alignment with neuromorphic hardware creates opportunities for energy-efficient learning [180], with dynamical memristors offering promising potential for brain-inspired systems [181]. Beyond conventional hardware, PC could enable "intelligence-in-a-dish" technology through adaptive computation in biophysical mediums [182, 183, 184]. Given the association of PC with cortical computation [144, 149, 185], it is inherently compatible with organoid intelligence [186]. Early spike-level implementations [187, 165] show promise for dynamic predictive processing. Organoid growth and decay could facilitate model selection [188], potentially improving generalization through natural evolution of neural structures.
Figure 8: Spiking neural coding formulation where $\mathbf{s}^{\ell}(t)$ represents spike nodes and $\mathbf{x}^{\ell}(t)$ are trace approximations. Solid arrows show forward synapses, dashed arrows show error transmission.
Spiking Predictive Coding.
A key challenge involves generalizing PC to spike-level processing. Biological neurons communicate through sparse action potentials creating information-rich spike trains [189], inspiring spiking neural networks that encode information through precise timing [190]. This enables energy-efficient neuromorphic hardware [191, 192, 193], unlike energy-intensive GPU-based networks. However, most PC formulations fail to account for sparse, discrete communication, reflecting the nature of PC as describing ensemble rather than individual neuron dynamics [131]. The spiking neural coding framework attempts PC implementation in spiking networks using spike-induced synaptic adjustments [194, 165]. Unlike standard spike-timing-dependent plasticity requiring hand-crafted layers [195], this framework uses flexible synaptic conductance models, membrane potentials, and trace mechanisms. Each layer estimates the trace activity of the following layer, passing discrepancies through evolved error feedback synapses (Fig. 8). The framework adapts to various spiking functions from leaky integrators to Hodgkin-Huxley models [196]. Recent implementations [197, 198, 152] operate in discriminative contexts but do not address weight transport, sharing feedback with forward synapses. For comprehensive coverage, see [199].
@@ 7 Looking Forward: Important Directions for Predictive Coding Research
As we have discussed in Sections 3 and 4, the performance of PC models can be quite below that of modern-day deep neural networks trained with backprop. The future of PC in machine learning strongly depends on our ability to address and fill this gap. Specifically, among the main directions for future work related to PC should be understanding the fundamental reasons behind this performance mismatch and using this acquired insight to develop and design novel PC schemas, mathematical frameworks, and computational models that work in large-scale settings, where current deep learning models excel.
@@
Future research will also need to focus on the study of optimization techniques that have proven useful and invaluable for variational inference, such as those that promote the inclusion of precision weighting parameters into the picture [206]. In the same spirit, recent progress on natural-gradient methods for gradient-based variational inference—ranging from stochastic NGVI with non-asymptotic guarantees, through Kronecker-factored curvature approximations, to the variational predictive natural gradient—offers curvature-aware updates that can be interpreted as learning layer-wise precisions online, and therefore provides an immediately transferable optimisation tool-set for scaling precision-weighted predictive coding [207, 208, 209]. Despite interesting developments obtained over the last several years, and despite the fact that precisions are of primary importance in simulations used in the neurosciences, these techniques have only been tested in small and medium scale settings [210, 211, 212]. More generally, the field of deep learning has enjoyed immense benefits from simple optimization tricks developed throughout the last decade, such as dropout [213], batch normalization [214], adaptive learning rates such as Adam [215] and RMSprop [216], and the introduction of the ReLU activation [217] and its variants. Without these techniques, emerging from the joint effort of thousands of researchers, the training of extremely overparametrized neural models would not lead to the results that we observe today. Interestingly, backpropagation itself exhibited many flaws and limitations at the beginning that precluded it scaling to high-dimensional data spaces: in the late 90s, kernel learning methods were dominant and generally found to be more effective than artificial neural networks, which were generally disregarded due the problem of vanishing gradients [218] as well as their strong requirement for heavy computational power. Most of these problems have now been addressed in over thirty years of research, again through a collective effort of thousands of people. Given the history of deep learning, one might ask: what are the dropout, batch norm, and Adam optimizer equivalents for PC? Addressing such a question would be of vital importance if the goal is to scale the applicability of PC, and will hopefully become a more prominent topic of interest of more research efforts going forward.
@@
Another probabilistic generative modeling research direction for PC is to incorporate the uncertainty about model (i.e., synaptic) parameters into the variational bound on marginal likelihood. That is, one can equip synaptic parameters with probability distributions (as opposed to using point estimates) [220]. This move would take PC closer to its Bayesian roots: current machine learning implementations of PC do not treat the model parameters as random variables and can therefore be regarded as an expectation maximization (EM) procedure, where the M-step ignores uncertainty about the parameters [221]. The benefit of treating parameters as random variables is that one can evaluate the model evidence required for structure learning [222, 223, 224, 225], i.e., one could place prediction errors on the synaptic weights and evaluate the resulting variational free energy (or marginal likelihood) for structure learning or, in a biological setting, morphogenesis [226, 227]. This research direction takes PC into the world of generalized filtering and dynamic causal modeling, in which parameters can be regarded as latent states that change very slowly [228, 229]. In the case of dynamic causal modeling (i.e., the variational inversion of state-space models under Gaussian assumptions), structure learning can be rendered particularly efficient through particular instances of Bayesian model selection, which provides a principled basis for comparing alternative generative hypotheses [230, 231, 232]. In predictive coding (PC), the same logic appears as Bayesian model reduction, where synaptic weights that receive little posterior evidence are pruned from a full (parent) network [233]. Looking ahead, a natural extension is to replace these fixed-dimension reductions with Bayesian non-parametric (BNP) priors—such as Dirichlet or Indian-buffet processes—that equip each PC layer with an unbounded supply of latent causes or synaptic motifs [231, 234, 235]. Thus in BNP–PC hybrid, precision-weighted prediction errors might simultaneously (i) adjust beliefs about existing components and (ii) signal when residual error warrants “breaking a new stick” (creating a component) or dropping a redundant one. Thus, BNP priors transform PC into a self-organising system that learns both its parameters and its effective dimensionality online, unifying synaptic growth and pruning within a single, neurally plausible free-energy minimisation scheme.
Crafting Control Systems.
@@
PC brings with it the promise of learning a powerful generative model that is continuously and iteratively refined as more sensory samples are gathered over time. This has led some early work to consider such a process as the basis for world models that drive modular, brain-motivated cognitive models, capable of combining perception and action in the context of playing video games [96, 129] and robotic control tasks [121, 119, 120], as well as large-scale cognitive architectures [236, 94, 126]. This has important implications beyond machine learning, particularly for the domains of cognitive science and cognitive neuroscience where a key pathway is to craft computational theories of mind and to examine their fit to human subject data on controlled psychological tasks [237] as well as their ability to generalize to cognitive functionality. Hence, a promising future direction is to design modular, increasingly complex systems made up of PC circuitry, that rely on single fundamental PC circuits and low-level dynamics.
@@
Using PC as a fundamental neural building block for control systems is still in its earliest stages, and although more effort will be needed for it to become viable for building robust computational theories of mind, it presents a worthwhile long-term direction for crafting simulated natural intelligence. This promise is made even stronger when considering that deep backprop-based networks are beginning to find use in the field of cognitive science [241]. Making developments along this direction would also crucially benefit the research in active inference [242, 92, 19, 18], given that agent systems that engage in epistemic foraging often center around the use of dynamic generative models and can even be viewed as simple control models.